OpenStudy (anonymous):

compare 6/9 and 10/12?

OpenStudy (texaschic101):

compare them ? well...6/9 = 2/3 and 10/12 = 5/6 2/3 = 4/6, so 10/12 is bigger. What else do you need ?

OpenStudy (anonymous):

That's it for now. Thanks

OpenStudy (anonymous):

Could you help me on this one? Solve 0.004 x= 52 @texaschic101

OpenStudy (texaschic101):

.004x = 52 (divide both sides by .004) .004x/.004 = 52/.004 x = 13,000 lets check it... .004(13,000) = 52 52 = 52 (correct) ANSWER : x = 13,000

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

Nope thank you! @texaschic101

OpenStudy (texaschic101):

anytime :) Glad I could help

OpenStudy (anonymous):

I'm sorry one more... The average size for the first five units was 20. The average size for the next 13 units was 77. What was the overall average? @texaschic101

OpenStudy (texaschic101):

overall average = (20 + 77)/2 = 97/2 = 48 1/2 or 48.5 I think....I am not 100% positive.